If del f =2xyz3 i + x2z3 j+ 3x2y2z2 k find the f(x,y,z) at f(1,-2,2)


1
Expert's answer
2021-03-16T08:11:27-0400

"f_x=2xyz^3,~f_y=x^2z^3,~f_z=3x^2yz^2,"

"f=x^2yz^3,"

"f(1,-2,2)=1^2\cdot(-2)\cdot2^3=-16."